# MVPatch: More Vivid Patch for Adversarial Camouflaged Attacks on Object Detectors in the Physical World

### v2

@@ IV Experiments
In this section, we briefly introduce the environment and dataset in the experiment. Additionally, we propose evaluation metrics for the experiment. In the digital environment, we generate meaningful and meaningless adversarial patches and compare them with similar patch attack approaches. In the physical environment, we compare the Attack Success Rate (ASR) of MVPatch with similar patch attack methods and vary the angles and distances between the patch and camera to investigate their impacts on the ASR.
@@ IV-A Implementation Details
IV-B Experimental Environment and Dataset
@@ IV-C Evaluation Metrics
mAP (mean Average Precision): The mAP is a widely used metric calculated by summing the average precision of models for each test sample and dividing it by the total number of test samples. It allows for comparisons of model performance across different datasets. The precision PREC is computed as the ratio of true positive samples divided by the sum of true positive and false positive samples. The recall REC is calculated as the ratio of true positive samples divided by the sum of true positive and false negative samples. Formally,
@@
NS (Naturalness Score): The NS is used to measure the similarity between the adversarial patch and the specified image, which is defined as
$\mathrm{NS}=$ $\lambda\left(\frac{\mathrm{CSS}-\widehat{\mathrm{CSS}}_{\mathrm{NG}}}{\widehat{\mathrm{CSS}}_{\mathrm{S}}-\widehat{\mathrm{CSS}}_{\mathrm{NG}}}+\frac{\mathrm{CSS}-\widehat{\mathrm{CSS}}_{\mathrm{NR}}}{\widehat{\mathrm{CSS}}_{\mathrm{S}}-\widehat{\mathrm{CSS}}_{\mathrm{NR}}}\right)$ (12)
@@
where CSS represents the cosine similarity score between the specified image and adversarial patch and $\widehat{\text{CSS}}$ is the cosine similarity score between the specified image and adversarial patch of NG and NR images. ED denotes the Euclidean distance. NG and NR are gray-scale and random noise. S is an abbreviation for the source image. A higher naturalness score indicates that MVPatch has a stronger similarity to the specified image.
TS (Transferability Score): The TS measures the transferable attack performance of the adversarial patch across diverse object detection models, which is defined as
@@
where $D_{i}$ denotes the mAP of object detection models detecting normal images while $\widehat{D}_{i}$ represents the mAP of object detection models detecting NG and NR images. NG and NR represent gray-scale and random noise. $||\cdot||_{2}$ is the L2 norm to constrain the distances between $D_{i}$ and $\widehat{D}_{i}$. The higher transferability score reveals that the adversarial patch has a higher transferable attack performance. We define
$\mathrm{ASR}=\frac{1}{|\mathcal{D}_{\mathrm{test}}|}\sum_{\mathcal{I}\in\mathcal{D}_{\mathrm{test}}}\mathbbm{1}\{f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)\},$ (14)
as an indicator of the attack success rate, where $f(\cdot)$ represents the consequence of the models’ detection. $\mathcal{D}_{\mathrm{test}}$ denotes a set of benign images from test datasets. The ASR equals 1 when $f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)$, and equals 0 otherwise.
@@ V Results
This section presents empirical evidence that showcases the effectiveness of the proposed MVPatch through comprehensive evaluations conducted in both digital and physical domains. In the digital domain, we divided the experimental objects into meaningful adversarial patches and meaningless adversarial patches. In the physical domain, we assessed the patch attack performance from various angles and distances by utilizing a camera to capture images as input sources. The experimental results demonstrate that our method outperforms other similar adversarial patch attack methods, both in the digital and physical domains. In the last, we show the visual consequence of the experiment in the digital domain and physical domain.
V-A Evaluation of Naturalness in the Digital World
To assess the naturalness of the adversarial patch, we utilize the YOLO series models as threat models. The natural factor, represented by the parameter $\lambda$, is adjusted within the range of [1,3] to generate a variety of adversarial patches. The naturalness score (NS) is then calculated to quantify the naturalness of these patches. The generated adversarial patches with different naturalness scores are displayed in Fig.3. As we move from left to right on the coordinate axis, the naturalness score gradually increases, indicating a higher level of naturalness.
@@ V-A Evaluation of Naturalness in the Digital World
Explanation for Experimental Consequence: $\lambda=[1,3]$ represents the adversarial patches generated by diverse natural factors. SRC represents the source image. NoiseGrey and NoiseRandom are matrices with values of 0.5 and random values, respectively. We deploy the mAP (mean Average Precision), CS (Cosine Similarity), and ED (Euclidean Distance) to measure the attack performance and the similarity of the adversarial patch with the source image.
@@
Analysis and Conclusion for Experimental Consequence: As shown in Table II, we vary $\lambda$ from 1 to 3, resulting in an increase in the score of naturalness from 35.59 to 81.01. Simultaneously, the mAP increases from 32.30% to 39.69%. However, a higher mAP generally corresponds to greater accuracy of the object detection models and a lower attack performance of adversarial patches. There is a tradeoff between naturalness and attack performance, so it is not possible to obtain an adversarial patch with both higher naturalness and attack performance. We identify a threshold for the score of naturalness that helps us obtain the optimal adversarial patch with both naturalness and attack performance. Through an analysis of 20 experimental participants, it is generally agreed that when the score of naturalness is above 75 ($\lambda=2.5$), the generated adversarial patch would be considered more natural. Furthermore, if we continue to increase the natural factor $\lambda$, the gain in naturalness recognizable to the human eye will decrease. Therefore, we define $\lambda=2.5$ as the parameter of the loss function in the following experiments. The images with different scores of naturalness for the adversarial patches are shown in Fig.3.
@@
${}^{(MeaningfulAdversarialPatches)}$hreat Models Transfer Attack Models Black Box(TS $\uparrow$)
@@
V-B Evaluation of Attack Transferability in the Digital domain
@@
To systematically investigate the attack performance of adversarial patches in the digital domain, we divide the experimental objects into two parts: meaningful adversarial patches and meaningless adversarial patches. Most research mainly focused on meaningless adversarial patches and few researchers investigated the meaningful and meaningless adversarial patches in the digital and physical domains.
V-B1 Evaluation of Attack Transferability of Meaningful Adversarial Patches in the Digital domain
@@
We utilize the pre-trained models including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny to generate diverse adversarial patches with meaningful content using the natural factor $\lambda=2.5$, as shown in Fig.4, 4, 4, 4, and 4. By combining these pre-trained models, we obtain ensemble attack adversarial patches, as shown in Fig.4 and 4. To compare the influence of person images and other images as adversarial patches for the object detection models, we use the sunflower image (Fig.4) as a control test for the person image (Fig.4). We input all the acquired adversarial patches into transfer attack models, including Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5, to calculate the transferability score and compare it with the Natural Patch[1]. The comparable experimental results are shown in Table III.
Explanation for Experimental Consequence: YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, YOLOv5, Faster R-CNN, and SSD are pre-trained object detection models. The ”Threat Models” column represents diverse adversarial patches generated by different models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and the Ensemble Model, which combines YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny. The ”original image” refers to an image without an adversarial attack.”Grey-scale Noise” and ”Random Noise” separately indicate an image with padding of a value of 0.5 and an image with random values. The Black Box (TS) measures the transferability of adversarial attacks, with the transferability score (TS) indicating superior performance with a higher score.
@@
Evaluation and Analysis for Experimental Consequence: Firstly, we compare the transferability score of MVPatch (Ours) to that of Natural Patch[1] on individual object detection models. MVPatch performs better than Natural Patch on most of the models in the experiment. For the YOLOv4 model, the TS of MVPatch is approximately 33% higher than the TS of Natural Patch. Similarly, the excellent attack performance of MVPatch is observed on the other four object detection models (YOLOv2, YOLOv3, YOLOv3-tiny, and YOLOv4-tiny). On the other hand, the attack performance of MVPatch on the ensemble model is surpassed by Natural Patch, with the TS of MVPatch on the ensemble model being approximately 16% higher than the TS of Natural Patch.
From the results of the meaningful adversarial patch experiment, it can be seen that the proposed algorithm for meaningful adversarial patches, MVPatch, achieves superior performance compared to the similar algorithm, Natural Patch, in terms of attack transferability.
V-B2 Evaluation of Attack Transferability of Meaningless Adversarial Patches in the Digital domain
We adjust the natural factor of $\lambda=0$ and utilize the joint pre-trained models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny, to generate meaningless adversarial patch MVPatch. In comparison, we consider AdvPatch[2], AdvTexture[3], and AdvCloak[32] as subjects for the experiment on meaningless adversarial patches. We use all the obtained adversarial patches as input for Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5, calculate the transferability scores, and compare them with those of AdvPatch [2], AdvTexture[3], and AdvCloak[32]. The comparable experimental results are presented in TableIV.
Explanation for Experimental Consequence: YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, YOLOv5, Faster R-CNN, and SSD are pre-trained object detection models. The ”Meaningless Adversarial Patches” column represents various meaningless adversarial patches generated by diverse attack algorithms. The Black Box (TS) measures the transferability of adversarial attacks, with a higher transferability score indicating better performance. TS stands for the transferability score.
@@ V-B2 Evaluation of Attack Transferability of Meaningless Adversarial Patches in the Digital domain
Evaluation and Analysis for Experimental Consequence: As illustrated in Table IV, the transferability score (TS) of MVPatch is approximately 40% higher than that of AdvCloak. Compared to AdvPatch and AdvTexture, the TS of MVPatch is approximately 20% higher. Moreover, when the transfer models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, YOLOv5, Faster R-CNN, and SSD, are attacked by MVPatch, their mean Average Precision (mAP) reduces by 10 to 20%, compared to the other three algorithms.
To sum up, in terms of meaningless adversarial patches, the proposed algorithm, MVPatch, exhibits higher attack transferability compared to similar algorithms.
V-C Evaluation of Attack Performance of Adversarial Patches in the Physical domain
@@
V-C1 Evaluation of Attack Performance of Diverse Adversarial Patches in the Physical Domain
The volunteers hold various adversarial patches, including AdvPatch, AdvTexture, a benign image, Natural Patch, and the proposed MVPatch, against their chests. The patches are positioned at an angle of {$0^{\circ},30^{\circ},60^{\circ},90^{\circ}$} and a distance of {1, 2, 3, 4} meters. We use an iPhone 11 to capture the images with the adversarial patches from diverse distances and angles in the physical world. In total, we collected 9681 images. To preserve the pixel details of the adversarial patches, we view them using an iPad Pro. The attack success rates and some evaluation metrics of different adversarial patches are shown in Table V, and sample patches are displayed in Fig.5.
Explanation for Experimental Consequence:The AdvPatch[2], and AdvTetxure[3] are both meaningless adversarial patches and the Natural Patches[1] and proposed MVPatches are meaningful patches. The Benign Image refers to an image without adversarial examples. ASR is short for attack success rate. Number of Images represents the number of experimental images to evaluate the attack performance. Stealthiness Attack means that the attack has a more naturalness score. Physical Attack denotes that the attack can be applied in the physical world. Transferable Attack represents that the attack has transferable attack ability. A lightweight model means that the model, that generates the adversarial patches, has low computational cost.

### v3

@@ IV Experiments
In this section, we provide a brief introduction to the environment and dataset used in the experiment. Additionally, we delineate the threat models considered in the experiment and introduce the evaluation metrics employed. In the digital environment, we generate both meaningful and meaningless adversarial patches and compare them with similar patch attack approaches. In the physical environment, we compare the Attack Success Rate (ASR) of MVPatch with similar patch attack methods and vary the angles and distances between the patch and the camera to investigate their impacts on the ASR.
@@ IV-A Implementation Details
IV-B Threat Models
IV-C Experimental Environment and Dataset
@@ IV-D Evaluation Metrics
mean Average Precision (mAP): The mAP is a widely used metric calculated by summing the average precision of models for each test sample and dividing it by the total number of test samples. It allows for comparisons of model performance across different datasets. The precision ‘PREC’ is computed as the ratio of true positive samples divided by the sum of true positive and false positive samples. The recall ‘REC’ is calculated as the ratio of true positive samples divided by the sum of true positive and false negative samples. Formally,
@@
Naturalness Score (NS): The NS is used to measure the similarity between the adversarial patch and the specified image, which is defined as
$\mathrm{NS}=$ $\lambda\left(\frac{\mathrm{CS}-\widehat{\mathrm{CS}}_{\mathrm{NG}}}{\widehat{\mathrm{CS}}_{\mathrm{S}}-\widehat{\mathrm{CS}}_{\mathrm{NG}}}+\frac{\mathrm{CS}-\widehat{\mathrm{CS}}_{\mathrm{NR}}}{\widehat{\mathrm{CS}}_{\mathrm{S}}-\widehat{\mathrm{CS}}_{\mathrm{NR}}}\right)$ (26)
@@
where CS denotes the cosine similarity score between the specified image S and the adversarial patch. $\widehat{\text{CS}}$ signifies the cosine similarity score between the specified image S and the adversarial patch of grayscale NG and random noise NR. Furthermore, ED represents the Euclidean distance. A higher naturalness score implies that the generated adversarial patch exhibits a stronger resemblance to the specified image.
Transferability Score (TS): The TS measures the transferable attack performance of the adversarial patch across diverse object detection models, which is defined as
@@
where $D_{i}$ denotes the mean Average Precision (mAP) of object detection models in identifying normal images, whereas $\widehat{D}i$ represents the mAP of object detection models in identifying NG (gray-scale) and NR (random noise) images. The symbols NG and NR respectively denote grayscale and random noise images. The term $||\cdot||_{2}$ refers to the L2 norm, which is employed to constrain the distances between $D_{i}$ and $\widehat{D}i$. A higher transferability score indicates that the adversarial patch demonstrates superior transferable attack performance.
The Attack Success Rate (ASR) is defined as follows:
$\mathrm{ASR}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{\mathcal{I}_{\mathrm{test}}}\mathbbm{1}\{f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)\},$ (28)
where $\mathcal{I}_{\mathrm{test}}$ denotes a set of benign images derived from test datasets, and $f(\cdot)$ represents the outcome of the models’ detection process. The ASR serves as an indicator of the attack success rate. It equals 1 if $f(\mathcal{I^{*},\theta})\neq f(\mathcal{I},\theta)$, and 0 otherwise.
@@ V Results
In this section, we present empirical evidence that underscores the efficacy of the proposed MVPatch through extensive evaluations conducted in both digital and physical domains. Initially, we illustrate the experimental results of various ensemble model strategies, demonstrating that our proposed ensemble model strategy exhibits superior performance compared to alternative approaches. Additionally, we evaluate naturalness by varying the natural factor $\lambda$. Within the digital domain, we categorize the experimental subjects into meaningful and meaningless adversarial patches. In the physical domain, we assess the performance of the patch attack from different angles and distances by employing a camera to capture images as input sources. The experimental outcomes reveal that our method surpasses other comparable adversarial patch attack methods in both digital and physical domains. Lastly, we present the visual outcomes of the experiments conducted in both the digital and physical domains.
V-A Evaluation of Transferability with Diverse Ensemble Model Strategies
Based on theoretical analysis discussed in Section III-C concerning the enhancement of transferability, the ensemble modeling technique can achieve superior adversarial transferability performance across multiple models when compared to a single model. There are, however, two primary approaches for leveraging the ensemble method: maximizing the outputs of neural networks and averaging the outputs of neural networks. We first investigate the utilization of two distinct loss functions for attacking object detection models, as outlined below:
@@ V-A Evaluation of Transferability with Diverse Ensemble Model Strategies
$\mathcal{L}_{\mathrm{OBJ_{1}}}=\frac{1}{M}\sum_{i=1}^{M}f_{i}(\hat{x}),$ (29)
$\mathcal{L}_{\mathrm{OBJ_{2}}}=\max\{f_{i}(\hat{x})\}^{M}_{i=1},$ (30)
where $f(\cdot)$ denotes the object confidence score of the $i^{th}$ object detection model in the ensemble of object detection models (YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny). $\hat{x}\in\mathcal{I^{*}}$ is the image with adversarial patch.
We compare the attack performance of diverse object confidence score loss functions on various object detection models. It is observed that the adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ outperforms the one generated by $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ in terms of attacking the object detection model. The adversarial patches generated through diverse loss functions are then applied to eight object detection models (Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models) to assess their attack performance and similarity to specified images. The adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ not only exhibits excellent attack capability (the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ is 32.20% and the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ is 51.80%) but also maintains a high level of naturalness. Consequently, we select $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ as our objective loss function.
@@
The results of the experiment are illustrated in Table I, where the Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models are included as threat models. The mean mAP represents the average value of the mean Average Precision (mAP) obtained from diverse models. CS corresponds to the Cosine Similarity and ED denotes the Euclidean Distance.
V-B Evaluation of Naturalness with Diverse Natural Factors
We assessed adversarial patches generated using YOLO models, adjusting the natural factor $\lambda$ within [1,3], and found that increasing $\lambda$ enhances naturalness, measured by the naturalness score (NS). The attack performance and similarity to the source image (SRC) were evaluated using mean Average Precision (mAP), Cosine Similarity (CS), and Euclidean Distance (ED). Additionally, NoiseGrey and NoiseRandom are matrices with values of 0.5 and random values, respectively.
As shown in Table II, varying $\lambda$ from 1 to 3 increases the naturalness score (NS) from 35.59 to 81.01. Simultaneously, the mean Average Precision (mAP) rises from 32.30% to 39.69%. However, a higher mAP generally indicates greater accuracy of the object detection models, leading to a lower attack performance of the adversarial patches. This highlights a tradeoff between naturalness and attack performance, making it challenging to achieve both high naturalness and high attack performance in a single adversarial patch.
Through analysis, we identify a threshold for naturalness that optimizes both aspects. Feedback from 20 experimental participants indicates that a naturalness score above 75 (achieved with $\lambda=2.5$) is generally perceived as more natural. Beyond this point, further increases in $\lambda$ result in diminishing gains in naturalness as perceived by the human eye. Therefore, we define $\lambda=2.5$ as the optimal parameter for the loss function in subsequent experiments. Figure 3 illustrates images of adversarial patches with different naturalness scores.
@@ V-B Evaluation of Naturalness with Diverse Natural Factors
(MeaningfulAdversarialPatches)Threat Models Transfer Attack Models Black Box(TS $\uparrow$)
@@
V-C Evaluation in the Digital Domain
@@
To systematically investigate the attack performance of adversarial patches, we categorize the experiments into two parts: meaningful and meaningless adversarial patches. Most research has primarily focused on meaningless patches, while the exploration of both meaningful and meaningless patches in digital and physical domains remains limited.
V-C1 Meaningful Adversarial Patches
@@
We utilize pre-trained models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny, to generate diverse adversarial patches with meaningful content using a natural factor of $\lambda=2.5$, as shown in Figs. 4 to 4. By combining these models, we create ensemble attack adversarial patches (Figs. 4 and 4). To compare the effects of person images and other images, we use a sunflower image as a control for the person image. All generated patches are tested against transfer attack models, including Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5, to calculate transferability scores (TS) and compare them with Natural Patch [1]. The results are presented in Table III. The ”Threat Models” column indicates adversarial patches generated by various models, while ”Grey-scale Noise” and ”Random Noise” refer to images padded with a value of 0.5 and random values, respectively. Higher TS indicates better transferability of the adversarial attacks.
@@
Analysis and Discussion: We compare the transferability score (TS) of MVPatch (Ours) to that of Natural Patch [1] on individual object detection models. MVPatch performs better than Natural Patch on most models in the experiment. For the YOLOv4 model, the TS of MVPatch is approximately 33% higher than that of Natural Patch. Similarly, MVPatch demonstrates excellent attack performance on other object detection models, including YOLOv2, YOLOv3, YOLOv3-tiny, and YOLOv4-tiny. Additionally, MVPatch surpasses Natural Patch in terms of attack performance on the ensemble model, with the TS of MVPatch being approximately 16% higher than that of Natural Patch. The results of the meaningful adversarial patch experiment clearly indicate that the proposed algorithm, MVPatch, achieves superior performance compared to the similar algorithm, Natural Patch, in terms of attack transferability.
V-C2 Meaningless Adversarial Patches
We set the natural factor to $\lambda=0$ and employ joint pre-trained models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny, to generate the meaningless adversarial patch. We compare this with AdvPatch [2], AdvTexture [3], and AdvCloak [33]. All generated patches are tested on Faster R-CNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 to calculate transferability scores (TS), which are then compared with the benchmarks. The results are summarized in Table IV. The ”Meaningless Adversarial Patches” column lists patches generated by various attack algorithms, while the Black Box (TS) indicates transferability, with higher scores reflecting superior performance.
@@ V-C2 Meaningless Adversarial Patches
Analysis and Discussion: As illustrated in Table IV, the transferability score (TS) of MVPatch is approximately 40% higher than that of AdvCloak. Compared to AdvPatch and AdvTexture, the TS of MVPatch is approximately 20% higher. Moreover, when the transfer models, including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, YOLOv5, Faster R-CNN, and SSD, are attacked by MVPatch, their mean Average Precision (mAP) reduces by 10 to 20%, compared to the other three algorithms. To sum up, in terms of meaningless adversarial patches, the proposed algorithm, MVPatch, exhibits higher attack transferability compared to similar algorithms.
V-D Evaluation in the Physical Domain
@@
V-D1 Comparision with Other Adversarial Patches
Table V presents the empirical results of various adversarial patches within the physical domain, including AdvPatch, AdvTexture, a benign image, Natural Patch, and the proposed MVPatch. Volunteers held these patches against their chests at angles of $0^{\circ},30^{\circ},60^{\circ},90^{\circ}$ and distances of 1, 2, 3, 4 meters. Using an iPhone 11, we captured a total of 9,681 images from various angles and distances, and reviewed them on an iPad Pro to ensure pixel detail preservation. The table evaluates key metrics such as Attack Success Rate (ASR) on the YOLOv2 model, number of images, stealthiness, physical applicability, transferability, and computational cost. AdvPatch and AdvTexture are classified as meaningless adversarial patches, while Natural Patch and MVPatch are meaningful. The benign image contains no adversarial features. ASR measures the attack success rate, ”Number of Images” indicates the experimental samples, ”Stealthiness” reflects naturalness, ”Physical Attack” confirms real-world applicability, ”Transferable Attack” indicates cross-domain effectiveness, and ”Lightweight Model” denotes low computational cost. The proposed MVPatch demonstrates superior performance across all evaluated metrics.